FOURIER TRANSFORM EMISSION SPECTROSCOPY OF NEW VISIBLE SYSTEMS OF NbN

\underline{R. S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ont., Canada N2L 3G1.

The emission spectrum of NbN has been reinvestigated in the 15000–35000 cm^{−1} region using a Fourier transform spectrometer and some new transitions have been observed in the visible region. The bands observed in the 18000– 19800 cm^{−1} region have been assigned as a new ^{3}Π − X ^{3}Δ transition. Three bands with R heads near 19463.8, 19659.0 and 19757.0 cm^{−1} have been assigned as the 0–0 bands of the ^{3}Π_{2} − X ^{3}Δ_{3}, ^{3}Π_{1} − X ^{3}Δ_{2} and ^{3}Π_{0^{±}} − X ^{3}Δ_{1} sub-bands of the new transition. To higher wavenumbers, a 0–0 band with origin near 25409.9 cm^{−1} has been assigned as a ΔΩ = 0 transition with X ^{3}Δ_{2} as the lower state. A rotational analysis of the 0–0 and 0–1 bands of these sub-bands has been carried out and spectroscopic constants have been extracted. Two additional bands with origins near 25518.7 and 25534.8 cm^{−1} are ΔΩ = 0 bands having X ^{3}Δ_{1} as the common lower state. Most of excited levels are affected by perturbations.